OpenStudy (anonymous):

x intercept of y=sqrt(8x-7) ???

jimthompson5910 (jim_thompson5910):

To find the x-intercept, plug in y = 0 and solve for x

jimthompson5910 (jim_thompson5910):

So y=sqrt(8x-7) 0=sqrt(8x-7) 0^2 = 8x - 7 0 = 8x - 7 7 = 8x 7/8 = x x = 7/8 which tells us that the x-intercept is (7/8, 0)

OpenStudy (anonymous):

Ohh, i see. i was thinking too hard.

jimthompson5910 (jim_thompson5910):

I'm glad it makes sense